4) Given a line of uniform charge of length 2.0 m and total charge +2.1 μC, what is the magnitude and direction of the electric field 1.0 m away from the line, perpendicular to the center of the line charge? here 1.0 m
Answer #1

charge per unit length

L = 2.1 / 2 = 10.5 uC/m

R = 1 m

electric field at required point

E = L / ( 2 pi* epsilon_not* R)

E = 10.5*10^-6 / ( 2 *3.14* 8.85*10^-12* 1)

E = 1.89*10^5 N/m

==========

do co. comment in case any doubt will reply for sure.. Goodluck